OpenStudy (anonymous):

what is the square root of 800?

OpenStudy (stormfire1):

~28.28

OpenStudy (anonymous):

28.2842712

OpenStudy (anonymous):

\[20\sqrt{2}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

\[\sqrt{800} = \sqrt{8} \cdot \sqrt{100} = \sqrt{4} \cdot \sqrt{2} \cdot \sqrt{100} = 2 \cdot 10 \cdot \sqrt{2} = 20\sqrt{2} \approx 28.28427125\]

OpenStudy (anonymous):

20 * 1.414 = 28.28